WRITE AN EQUATION OF THE LINE THAT CONTAINS THE GIVEN POINTS.
A. (3,-3), (-3,1)
B. (3,5), (5,3)

Answer by benni1013(206)

(3,-3), (-3,1)
First you need to find the slope:
change in y over change in x. Make sure you back to the same ordered pair.
-3-1/3-(-3)==>-4/6==>-2/3
We have a slope of -2/3. Now plug in one of the given points into the slope-intercept form. I'll pick (-3,1).
y=mx+b==>1=3(-2/3)+b==>1=-2+b==>b=3
Plug back into slope-intercept form.
y=-2/3x+3